Cavity alignment was not recoverable. I most likely spent way too much time trying to recover instead of just starting from scratch. This is frustrating for future first contact cleaning if this cleaning procedure does improve the cavity finesse. The proccess I tried was the following:
1. Assume input alignment was unchanged, place both input couplers back on the cavity. Adjust input couplers to align both light paths back on their refl PDs. With how small the diode is on the high BW REFL PDs this should ensure 2 mirrors in their original position.
2. Piezo mirror (775 output coupler) is static and cannot be changed. Nothing to be done there.
3. Scan around on third adjustable mirror and look for any second pass beams on the camera.
It seems to me like this should work, but this did not yield anything.I believe the piezo mirror is at a drastically different angle, which is strange because when I was doing piezo science before, changing out the piezos did not drastically change the intercavity alignment. I started over by changing the input alignment of the 1550 light and adjusting the cavity mirrors to match the 3D printed alignment tools. Doing this instead I found cavity flashes within an hour. Realigned the REFL and TRANS light onto the cameras/PDs and fine tuned alignment. OFC1 1550 light is now aligned. With the newly built tank circuit the error signal on the 1550 path looks very clean. I believe the frequency of the modulations is not quite correct and can be looked into further. I have taken ringdown measurements to see if cleaning was successful, but am getting finesse numbers greater than the specs of the mirrors. I think this may be due to the TRANS pd being saturated? Going to retry. Also need to double check the 1550 TRANS PD is fast enough, although for the 1550 light the decay time should be on the order of 3300/pi/c*2.4 = 8e-6 s or 100's of kHz. I think all the PDs we have should be this fast.
Additionally, i took off the 775 input coupler and tried to match the input alignment. As built, the 775 light first sees the piezo mirror after the input coupler. For any given input alignment the piezo mirror is putting the beam into the metal casing of the cavity. Need to explore this still.
Wow, that's awesome to hear. Good optics and good procedures.
Update:
Retook the measurement, making sure to move the laser frequency away from equilibrium enough via the output offset, so that when I break lock no new light is resonant and I am purely measuring the light exiting the cavity. Additionally reduced input power slightly so that the PD is not saturated. Did two measurements in transmission of the cavity and find finesse measurements very close to the specs of the mirror. I think more analysis is needed (align 775 and do similar measurements) but we may be able to conclude that the mirrors were simply dirty coming from the manufacturer. Plot attached. It seems like this may be worth doing for the second cavity.